# Analysis of a multi-target linear shrinkage covariance estimator

### v1

@@ Abstract
Keywords: Multi-target, linear shrinkage, covariance estimation, general asymptotics
@@ 1 Introduction and related work
The covariance matrix plays a major role in numerous machine learning algorithms and statistics. Just to cite a few, array signal processing [1], [2], generalized method of moments estimators [3], financial portfolio management [4], [5] and forecasting [6], functional genomics [7]. However, those algorithms are designed to use the population covariance matrix, which is often unaccessible. The sample covariance matrix severely fails in many applications: for instance, for Markowitz portfolio management it does often not beat a naive uniform diversification among the assets [8].
In the context of Kolmogorov asymptotics, where the ratio of the dimension $p_{n}$ and the number of samples $n$ tends to a finite positive constant $\frac{p_{n}}{n}\rightarrow c>0$, the sample covariance eigenvalue spectrum is spread: high eigenvalues tends at being too high, and low ones, too low [9], [10], [11].
@@
More formally, we denote by $S$ the sample covariance, and we consider $N\in\mathbb{N}$ targets, denoted $(T^{(i)})_{i=1}^{N}$, which are constant matrices of the same size as $S$. Usually, when there is only one target, $N=1$, we chose $T^{(1)}=I$ the identity matrix [13]. With $N=2$ targets, a standard example is the following. Consider the a diagonal block population covariance $\Sigma=\begin{pmatrix}\Sigma_{1}&(0)\\ (0)&\Sigma_{2}\end{pmatrix}.$ Even if $\Sigma_{1}$ and $\Sigma_{2}$ are unknown, we can leverage our knowledge of the block structure, and chose the following targets:
$T^{(1)}=\begin{pmatrix}I_{1}&(0)\\ (0)&(0)\end{pmatrix}\text{ and }T^{(2)}=\begin{pmatrix}(0)&(0)\\ (0)&I_{2}\end{pmatrix}.$ (1)
@@
$\hat{\Sigma}=c_{0}S+\sum_{i=1}^{N}c_{i}T^{(i)}\text{, where }(c_{0},...,c_{N})\in\mathbb{R}^{N}.$ (2)
@@
However, the performance of those algorithms highly depends on the choice of the target and its similarity to the population covariance $\Sigma$. Thus, there is no general ideal choice of target fitting all covariances $\Sigma$. The use of several targets instead of one can lead to superior classification performance [17, 18]. The methods to compute multi-target linear shrinkage estimators are less numerous. With only two targets, an analytic estimator is given in [17]. The general case for any number of targets an estimator was proposed for Gaussian distributions in [19]. In the distribution-free case, an implicit analytic estimator was proposed in [18], retrieved from solving a strictly convex problem. Nevertheless, to the best of our knowledge, there is no theoretical analysis of any of those algorithms comparable to the quadratic convergences we have in single-target shrinkage.
@@
For easier reading, technical proofs of the theoretical results are derived in the appendix, we will only give the idea of proof for intuition in the main corpus.
@@ 2 Notation, definitions and hypotheses
W define the sample covariance in the setting of known and unknown mean, and the set of targets.
@@
$S_{n}=\tilde{X}_{n}\tilde{X}_{n}^{T}/(n-1),\text{ with }(\tilde{X}_{n})_{ik}=(X_{n})_{ik}-\frac{1}{n}\sum_{k^{\prime}=1}^{n}(X_{n})_{ik^{\prime}}.$ (4)
@@
The “orthogonal and normalized” condition is not really a restriction in practice. Indeed, we can built any linearly independent family $T_{n}^{(k)}$ of symmetric matrices. Then, we can use the Gram-Schmidt algorithm [20] on this family to make it ortho-normalized, without modifying $\text{Span}\left(\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}\right)$, so without changing the space of multi-target estimators of the form $\hat{\Sigma}=c_{0}S+\sum_{i=1}^{N}c_{i}T^{(i)}\text{, where }(c_{0},...,c_{N})\in\mathbb{R}^{N}$.
@@ 3 Oracle and proposed estimator
Remark 1.
Given a sample covariance $S_{n}$, we are looking for a minimizer of the 2-norm in the scope of linear combinations of $S_{n}$ and $T_{n}^{(k)}$.
@@
The minimizer $c^{*}\in\mathbb{R}^{N_{n}+1}$, solve the following equation:
$A_{n}c^{*}=b_{n},\text{ with }$ $A_{n}=\begin{pmatrix}\lVert S_{n}\rVert_{n}^{2}&\langle S_{n},T_{n}^{(1)}\rangle_{n}&\ldots&\langle S_{n},T_{n}^{(N_{n})}\rangle\\ \langle S_{n},T_{n}^{(1)}\rangle_{n}&1&&(0)\\ \vdots&&\ddots&\\ \langle S_{n},T_{n}^{(N_{n})}\rangle&(0)&&1\end{pmatrix},$ (6)
We have:
$\det(A_{n})=\lVert S_{n}\rVert_{n}^{2}-\sum_{k}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2},$ (7)
$\text{com}(A_{n})^{T}=\begin{pmatrix}1&-\langle S_{n},T_{n}^{(1)}\rangle_{n}&\ldots&-\langle S_{n},T_{n}^{(N_{n})}\rangle\\ -\langle S_{n},T_{n}^{(1)}\rangle_{n}&&&\\ \vdots&&(B_{n})&\\ -\langle S_{n},T_{n}^{(N_{n})}\rangle&&&\end{pmatrix},$
$\text{with }B_{n}=\left(\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}\right)I_{N_{n}}+QQ^{T},Q=\left[\langle S_{n},T_{n}^{(i)}\rangle_{n}\right]_{i=1,...,N_{n}}^{T}.$
Remark 2.
Previous derivations lead to the following oracle estimators.
Definition 3 (Oracle estimators).
The oracle estimators solving the minimization problem 5 are, for $i\in\llbracket 1,N_{n}\rrbracket$:
$c_{0}^{*}=\frac{1}{\det(A_{n})}\left(\langle S_{n},\Sigma_{n}\rangle_{n}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}\langle\Sigma_{n},T_{n}^{(k)}\rangle_{n}\right),c_{i}^{*}=\langle\Sigma_{n},T_{n}^{(i)}\rangle_{n}-c_{0}^{*}\langle S_{n},T_{n}^{(i)}\rangle_{n}.$ (8)
And we define the oracle estimator of covariance $\Sigma_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}^{*}S_{n}+\sum_{k=1}^{N_{n}}c_{k}^{*}T_{n}^{(k)}\right)$, where $\mathcal{P}_{\mathcal{S}_{n}^{+}}$ is the orthogonal projection on the cone of symmetric positive matrices.
The orthogonal projection $\mathcal{P}_{\mathcal{S}_{n}^{+}}$ at the end does not strictly follows the definition of linear shrinkage. However, as $\Sigma_{n}$ is symmetric positive, the projection step decreases the distance $\lVert\Sigma_{n}^{*}-\Sigma_{n}\rVert_{n}$. This step is not necessary for the convergence results to hold, but it is a free improvement we recommend.
@@
$\hat{\mathbb{V}}(S_{n})=\frac{1}{n(n-1)}\sum_{k=1}^{n}\left\lVert X_{\cdot k}^{n}\left(X_{\cdot k}^{n}\right)^{T}-S_{n}\right\rVert_{n}^{2}.$ (9)
@@
$\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})=\frac{1}{n(n-1)}\sum_{k=1}^{n}\left\langle X_{\cdot k}^{n}\left(X_{\cdot k}^{n}\right)^{T}-S_{n},T_{n}^{(i)}\right\rangle_{n}^{2}.$ (10)

### v2

@@ Abstract
keywords:
linear shrinkage, covariance estimation, high dimension, general asymptotics
Acknowledgments
We want to give special thanks to Alexandre Miot (Core ML Lab, Société Générale Corporate and Investment Banking) and Gabriel Turinici (CEREMADE, Université Dauphine-PSL) for their advices all along the work.
@@ 1 Introduction and related work
The covariance matrix plays a major role in numerous machine learning algorithms and statistics. Just to cite a few, array signal processing [1, 9], generalized method of moments estimators [10], financial portfolio management [18, 14] and forecasting [22], functional genomics [20]. However, those algorithms are designed to use the population covariance matrix, which is often unaccessible. The sample covariance matrix severely fails in many applications: for instance, for Markowitz portfolio management it does often not beat a naive uniform diversification among the assets [4].
In the context of Kolmogorov asymptotics, where the ratio of the dimension $p_{n}$ and the number of samples $n$ tends to a finite positive constant $\frac{p_{n}}{n}\rightarrow c>0$, the sample covariance eigenvalue spectrum is spread: high eigenvalues tends at being too high, and low ones, too low [5, 23, 6].
@@
More formally, we denote by $S$ the sample covariance, and we consider $N\in\mathbb{N}$ targets, denoted $(T^{(i)})_{i=1}^{N}$, which are constant matrices of the same size as $S$. Usually, when there is only one target, $N=1$, we chose $T^{(1)}=I$ the identity matrix [15]. With $N=2$ targets, a standard example is the following. Consider a diagonal block population covariance $\Sigma=\begin{pmatrix}\Sigma_{1}&(0)\\ (0)&\Sigma_{2}\end{pmatrix}.$ The method does not require $\Sigma$ to be block-diagonal, this example uses it only for intuition and explainatory reasons. Even if $\Sigma_{1}$ and $\Sigma_{2}$ are unknown, we can leverage our knowledge of the block structure, and chose the following targets:
$T^{(1)}=\begin{pmatrix}I_{1}&(0)\\ (0)&(0)\end{pmatrix}\text{ and }T^{(2)}=\begin{pmatrix}(0)&(0)\\ (0)&I_{2}\end{pmatrix}.$
@@
$\hat{\Sigma}=c_{0}S+\sum_{i=1}^{N}c_{i}T^{(i)}\text{, where }(c_{0},...,c_{N})\in\mathbb{R}^{N}.$
@@
However, the performance of those algorithms highly depends on the choice of the target and its similarity to the population covariance $\Sigma$. Thus, there is no general ideal choice of target fitting all covariances $\Sigma$. The use of several targets instead of one can lead to superior classification performance [11, 13]. The methods to compute multi-target linear shrinkage estimators are less numerous. With only two targets, an analytic estimator is given in [11]. The general case for any number of targets an estimator was proposed for Gaussian distributions in [7]. In the distribution-free case, an implicit analytic estimator was proposed in [13], retrieved from solving a strictly convex problem. To the best of our knowledge, there is no theoretical analysis of those algorithms comparable to the quadratic convergences we have in single-target shrinkage.
@@
For easier reading, technical proofs of the theoretical results are derived in the appendix, we will only give the idea of proof for intuition in the main corpus. The code and data for the experiment is available on our GitHub.
@@ 2 Notation, definitions and hypotheses
We define the sample covariance in the setting of known and unknown mean, and the set of targets.
@@
$S_{n}=\tilde{X}_{n}\tilde{X}_{n}^{T}/(n-1),$ (2)
with $(\tilde{X}_{n})_{ik}=(X_{n})_{ik}-\frac{1}{n}\sum_{k^{\prime}=1}^{n}(X_{n})_{ik^{\prime}}$.
@@
The “orthogonal and normalized” condition is not really a restriction in practice. Indeed, we can build any linearly independent family $T_{n}^{(k)}$ of symmetric matrices. Then, we can use the Gram-Schmidt algorithm [8] on this family to make it ortho-normalized, without modifying $\text{Span}\left(\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}\right)$, so without changing the space of multi-target estimators of the form $\hat{\Sigma}=c_{0}S+\sum_{i=1}^{N}c_{i}T^{(i)}\text{, where }(c_{0},...,c_{N})\in\mathbb{R}^{N}$.
@@ 3 Oracle and proposed estimator
Given a sample covariance $S_{n}$, we are looking for a minimizer of the 2-norm in the scope of linear combinations of $S_{n}$ and $T_{n}^{(k)}$, i.e. the orthogonal projection of the population covariance onto the subspace spanned by $\{S_{n},T_{n}^{(1)},...,T_{n}^{(N_{n})}\}$.
@@
The idea is to solve this minimization problem, construct an oracle covariance estimator with it, and then estimate its components and propose a bona fide covariance estimator. Firstly, let us see the oracle covariance estimator $\Sigma_{n}^{*}$ that emerges.
Proposition 1 (Oracle estimator).
The oracle coefficients solving the minimization problem 3 are, for $i\in\llbracket 1,N_{n}\rrbracket$, with $\langle\cdot\rangle:=\langle\cdot\rangle_{n}$:
$c_{0}^{*}=\frac{1}{d_{n}^{2}}\left(\langle S_{n},\Sigma_{n}\rangle-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle\langle\Sigma_{n},T_{n}^{(k)}\rangle\right),$ (4)
$c_{i}^{*}=\langle\Sigma_{n},T_{n}^{(i)}\rangle-c_{0}^{*}\langle S_{n},T_{n}^{(i)}\rangle,$
$d_{n}^{2}=\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}.$
Thus, we define the oracle covariance estimator $\Sigma_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}^{*}S_{n}+\sum_{k=1}^{N_{n}}c_{k}^{*}T_{n}^{(k)}\right)$, where $\mathcal{P}_{\mathcal{S}_{n}^{+}}$ is the orthogonal projection on the cone of positive semi-definite matrices.
Remark that even the formulas tend to look rough, the oracle estimator is none else but the orthogonal projection of the population covariance $\Sigma_{n}$ onto the subspace spanned by $\{S_{n},T_{n}^{(1)},...,T_{n}^{(N_{n})}\}$, followed by projection onto the cone of positive semi-definite matrices $\mathcal{S}_{n}^{+}$.
This orthogonal projection $\mathcal{P}_{\mathcal{S}_{n}^{+}}$ does not strictly follow the definition of linear shrinkage. However, as $\Sigma_{n}$ is symmetric positive, the projection step decreases the distance $\lVert\Sigma_{n}^{*}-\Sigma_{n}\rVert_{n}$. This step is not necessary for the convergence results to hold, but it is a free improvement we recommend.
@@
$\hat{\mathbb{V}}(S_{n})=\frac{1}{n(n-1)}\sum_{k=1}^{n}\left\lVert X_{\cdot k}^{n}\left(X_{\cdot k}^{n}\right)^{T}-S_{n}\right\rVert_{n}^{2}.$
@@
$\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})=\sum_{k=1}^{n}\frac{\left\langle X_{\cdot k}^{n}\left(X_{\cdot k}^{n}\right)^{T}-S_{n},T_{n}^{(i)}\right\rangle_{n}^{2}}{n(n-1)}.$
For the unknown mean setting, closed-form formulas are derived for $\hat{\mathbb{V}}(S_{n})$ and $\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$, and are detailed in the Appendix.
Theorem 1 (Variance estimators).
Denoting $\mathbb{V}(S_{n})=\mathbb{E}[\lVert S_{n}-\Sigma_{n}\rVert_{n}^{2}]$ and $\mathbb{V}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)=\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(i)}\rangle_{n}^{2}\right]$, under assumptions 1 to 4, both where the mean is known or unknown for all $i\in\llbracket 1,N_{n}\rrbracket$, $\hat{\mathbb{V}}(S_{n})$ and $\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ are unbiased estimators of $\mathbb{V}(S_{n})$ and $\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ respectively. Moreover, $\mathbb{E}\left[\left(\hat{\mathbb{V}}(S_{n})-\mathbb{V}(S_{n})\right)^{2}\right]\longrightarrow 0$ and $\mathbb{E}\left[\left(\sum_{i=1}^{N_{n}}\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})-\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)^{2}\right]\longrightarrow 0$.
The idea of proof of this theorem resides in carefully developing, counting null terms and using the assumptions 3 and 4 through Cauchy-Schwarz and Jensen inequalities.
The previous results gave us bona fide estimators - that behave well asyptotically - for each quantity necessary to compute $\Sigma_{n}^{*}$. Then, we can define our bona fide covariance estimator $S_{n}^{*}$. As we carefully chose estimators that behave well asymptotically, we prove that $S_{n}^{*}$ converges to $\Sigma_{n}^{*}$ in quadratic mean.
Definition 4 (Bona fide estimator).
The bona fide estimators $c$ for the minimization problem 3 are defined, for $i\in\llbracket 1,N_{n}\rrbracket$, as:
$c_{0}=\min\left((\hat{c}_{0})_{+},1\right),c_{i}=(1-c_{0})\langle S_{n},T_{n}^{(i)}\rangle_{n},$ (5)
@@
with $\text{with }\hat{c}_{0}=$ $\frac{1}{d_{n}^{2}}\Bigg{(}\lVert S_{n}\rVert_{n}^{2}-\hat{\mathbb{V}}(S_{n})-\sum_{k=1}^{N_{n}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}-\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)\right)\Bigg{)},$
$d_{n}^{2}=$ $\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2},$
The bona fide MTSE is defined as:
$S_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}S_{n}+(1-c_{0})\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}\right).$
The implementation in Python is available on our GitHub. The truncation of $c_{0}$ between 00 and $1$ theoretically does not change the asymptotic behavior of the estimator. Geometrically, it ensures that $S_{n}^{*}$ is a convex combination between $S_{n}$ and the matrix $\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}$. Empirically, it improves the robustness of the estimator, particularly in small sample size.
Theorem 2 (Loss convergence).
The idea of the proof is to prove the convergence on the estimators without the projection step onto the cone of positive matrices, and use the Lipschitz property of projection to conclude. The initial convergence relies on the Lemma A.1 in [15].
4 Experimental results
This section illustrates the practical consequences of the choice of targets on the estimation, the robustness of the estimator to heavy tails, and an application to real non-stationary time series. The experiments were computed on a Lenovo laptop, Intel i5, 32 GB Ram. For reproducibility, the code and data are available on our GitHub, and the settings of the experiments are detailed in the Appendix. The performance is measured with the Percentage Relative Improvement in Average Loss (PRIAL):
It is to be noted that the complexity of the implementation can be reduced to $O(N_{n}p_{n}^{3}+N_{n}np_{n})$. The details are in the Appendix.
4.1 Target set impact: Usefulness of a target
We fix the dimension at $p=50$, and $n=25$, and we study the impact of the number of targets $N_{n}$. The population covariance $\Sigma$ is designed to clearly exhibit a choice of useful targets, and compare the role of targets improve significantly the oracle PRIAL-, to those which give less. $\Sigma$ is made as follows - a formal definition is given in the Appendix -, where blocks $B_{i}$ are drawn in a Wishart distribution:
$\Sigma=\begin{pmatrix}B_{1}&&(0)\\ &\ddots&\\ (0)&&B_{10}\end{pmatrix}.$ (7)
We consider two sets of targets: one set of “aligned” targets $\{T_{1}^{(i)}\}$, and one set of “misaligned” ones $\{T_{2}^{(i)}\}$. The first one is built as following: the first target $T_{1}^{(1)}=I$, to match Ledoit-Wolf STSE [15], and for $i\in\llbracket 2,10\rrbracket$, the target $T_{1}^{(i)}$ is block-diagonal; where $\Sigma$ has the block $B_{i}$, $T_{1}^{(i)}$ has an Identity block:
$T_{1}^{(i)}=\begin{pmatrix}(0)&\ldots&(0)\\ \vdots&(\text{block }i:)I_{5}&\vdots\\ (0)&\ldots&(0)\end{pmatrix}$ (8)
For the second set of targets $\{T_{2}^{(i)}\}$, we break the perfect alignment between $T_{1}^{(i)}$ and the block $B_{i}$. For that, we permute the rows and columns: column $j$ in $\{T_{1}^{(i)}\}$ becomes column $j+2$ in $\{T_{2}^{(i)}\}$ and so on.
We consider the PRIAL of the MTSE (MTSE_1 and MTSE_2 using respectively $\{T_{1}^{(i)}\}$ and $\{T_{2}^{(i)}\}$), the oracle ones, and the Ledoit-Wolf STSE (LW), in function of the number $k$ of targets used, between $k=1$, which is equivalent to the STSE, and $k=10$ where the sets of targets are used completely. The results, with $50000$ draws and Student distribution ($\nu=9$ d.o.f.), are shown in Figure 1. The maximum standard deviation is $4\times 10^{-4}$, error bars were removed.
Figure 1: PRIAL of the MTSE, and the oracle, with target sets 1 and 2, and the Ledoit-Wolf STSE as reference.
Figure 2: PRIAL difference between the oracle MTSE and Ledoit Wolf estimator (MTSE_o - LW), and bona fide MTSE (MTSE - LW).
It confirms the intuition about a useful target: the “aligned” set $(T_{1})^{(i)}$ presents a significant better PRIAL than the second misaligned one $(T_{2})^{(i)}$. This comprehension of what makes a target useful is of utter importance to design the target set in a practical use case. The second conclusion is that even if the misaligned targets are not optimal, they still present a better PRIAL than the STSE using only one target.
4.2 Target set impact: Adding useless targets to a good target set
The first target is the identity matrix $I$, to match with Ledoit-Wolf STSE. For $i\in\llbracket 2,10\rrbracket$, $T^{(i)}$ is designed according to Equation 14. For target $i\geq 11$, the target is drawn from a Wishart distribution $\mathcal{W}_{p}(I_{p},p)$. The latter are supposed to be “useless” regarding the structure of $\Sigma$.
We ran the experiment for $5000$ times, the maximum standard deviation is $7\times 10^{-4}$. We plot the differences of PRIAL between the MTSE and the Ledoit-Wolf STSE, the same for the MTSE oracle. The result can be seen in figure 2.
Figure 3: Convergence of $\lVert S_{n}^{*}-\Sigma_{n}^{*}\rVert_{n}^{2}$ for different values of $\nu$.
4.3 Heavy-tails robustness
In the case of Ledoit-Wolf STSE, Ledoit and Wolf empirically argue that Assumption 2 about finite eighth moments is in practice not necessary, only finite fourth moments would suffice [15]. Thus, we study the convergence of $S_{n}^{*}$ towards $\Sigma_{n}^{*}$ in a setting of heavy tails. For that, we consider multivariate t-Student distributions with $\nu>2$ degrees of freedom, for different values of $\nu\in\{20000,8,4,3,2.5\}$. The choice of $\Sigma_{n}$ and targets is detailed in the Appendix.
This experiment, which results are in Figure 3, strengthens the hypothesis of Ledoit and Wolf [15] saying that bounded fourth moments is experimentally enough, and $\nu=3$ is what we found the be the frontier for the convergence to hold
4.4 Application to non-stationary time series
In order to give the practitioner a real example on an application, we consider a standard portfolio management problem in finance.
We use daily log-returns of $323$ assets that stayed in the S&P500 between 2010 and 2022  [^1] Data were retrieved from YahooFinance using yfinance package, and available with the code on GitHub.. We recall that the log-return $r_{t}$ at day $t$ of a price $s_{t}$ is defined as the log-increment: $r_{t}=\log(s_{t})-\log(s_{t-1})$. So we have a $323$-dimensional time series, with one vector per day.
Those $323$ assets are classified in $11$ industrial sectors, known as the GICS sectors. We expect to take advantage of it, using those sectors to build $11$ targets.
To evaluate the performance of different covariance estimators, we used the General Minimum Variance - GMV - portfolio problem, a standard method to compare covariance estimators. The idea of GMV is to determine at the beginning of each month $T$ a weight vector $w_{T}\in\mathbb{R}^{323}$, and invest $w_{i,T}$ on the asset $i$ for one month: this is our portfolio.
At each day $t$ of the month $T$, around $20$ business days, we collect the log-returns of our portfolio $R_{t}=w_{T}^{T}r_{t}$, and at the end of the month we store the empirical variance of our portfolio during the month: $v_{T}=\mathbb{V}[\{R_{t}\}_{t=T}^{T+20}]$. The weight vector $w_{T}$ is designed to minimize this variance $v_{T}$, under constraint that $w_{T}^{\top}1=1$. There is closed-form formula, where $\Sigma$ is the covariance matrix of the random vector $r_{t}$: $w_{T}=\frac{\Sigma^{-1}1}{1^{\top}\Sigma^{-1}1}$. The pseudo-inverse is used when not invertible.
The experiment goes as following: at the beginning of each month $T$, we estimate $\hat{\Sigma}$ on the last $K$ months, i.e roughly $20K$ data points, and we compute $\hat{w}_{T}=\frac{\hat{\Sigma}^{-1}1}{1^{T}\hat{\Sigma}^{-1}1}$. At the end of the month we store the empirical variance $\hat{v}_{T}$. We do this procedure over $135$ months, between 2010 and 2022, and we compare the Cumulative Variance $\sum_{T=1}^{135}v_{T}$, the results are shown in the Table 1. The better the estimator $\hat{\Sigma}$, the lower is the Cumulative Variance.
The number $K$ of months used to fit the covariance estimators varies between $3$ and $15$. Along with sample covariance $S$, we compare:
  • MTSE_L defined in [13], another multi-target estimator. The code being unavailable (as remarked in [7], Section 6), we implemented the algorithm specifically for our set of 11 GICS targets, using the fact that they are diagonal to get rid of the computation of $E_{q}\otimes E_{q}$ (Kronecker product) in [13] of size $327^{2}\times 327^{2}$, which would be unusable in practice,
  • MTSE_G defined in [7], another multi-target estimator. In this setting, we cannot use our 11 GICS targets (as they are not invertible), so We used the nine targets proposed in their work,
  • two STSE: LWO, Ledoit-Wolf estimator for unknown mean (see [19]), and OAS (see [3]),
  • two non-linear shrinkage: ANS (see [17]), and QIS (see [16]).
The MTSE makes the difference in this difficult and non-stationnary task, as it consistently has a lower Cumulative Variance than the other considered algorithms.
Estimator $3$ months $4$ months $6$ months $9$ months $12$ months $15$ months
MTSE $\mathbf{7.14\times 10^{-3}}$ $\mathbf{7.13\times 10^{-3}}$ $\mathbf{7.29\times 10^{-3}}$ $\mathbf{7.09\times 10^{-3}}$ $\mathbf{7.62\times 10^{-3}}$ $\mathbf{7.52\times 10^{-3}}$
MTSE_L $7.71\times 10^{-3}$ $7.76\times 10^{-3}$ $8.41\times 10^{-3}$ $8.28\times 10^{-3}$ $9.24\times 10^{-3}$ $8.66\times 10^{-3}$
MTSE_G $15.29\times 10^{-3}$ $15.28\times 10^{-3}$ $14.70\times 10^{-3}$ $14.39\times 10^{-3}$ $14.30\times 10^{-3}$ $14.19\times 10^{-3}$
LWO $7.36\times 10^{-3}$ $7.22\times 10^{-3}$ $7.40\times 10^{-3}$ $7.25\times 10^{-3}$ $7.91\times 10^{-3}$ $7.65\times 10^{-3}$
OAS $7.41\times 10^{-3}$ $7.34\times 10^{-3}$ $7.77\times 10^{-3}$ $7.90\times 10^{-3}$ $8.99\times 10^{-3}$ $8.90\times 10^{-3}$
ANS $7.39\times 10^{-3}$ $7.34\times 10^{-3}$ $7.91\times 10^{-3}$ $8.42\times 10^{-3}$ $10.95\times 10^{-3}$ $12.08\times 10^{-3}$
QIS $1.38\times 10^{0}$ $33.22\times 10^{-3}$ $8.80\times 10^{-3}$ $10.45\times 10^{-3}$ $17.56\times 10^{-3}$ $53.60\times 10^{-3}$
S $9.15\times 10^{-3}$ $8.64\times 10^{-3}$ $9.38\times 10^{-3}$ $11.54\times 10^{-3}$ $16.40\times 10^{-3}$ $45.22\times 10^{-3}$
5 Limitations
The strongest assumptions made in the theoretical analysis are that we have i.i.d. samples and bounded eighth moments. In practice, the i.i.d. assumption is hard to assess, particularly in time series analysis, as we saw in the financial dataset. The requirement of finite eighth moments can be a limitation, even if we saw that experimentally the convergence holds for heavier tails.
As in single-target estimation, the central point for performance in a task is a meaningful choice of targets. Some works address the topic in the literature, mostly for Single-Target estimators, see [14, 24] for example. For multi-target estimators, [13] proposes a set of four targets in Section III.B, and [7] proposes a set of nine targets (section 2.4). However, a general method to design them and choose their number still lack in the literature to the best of our knowledge.
6 Conclusion
We derived the oracle estimator solving the multi-target linear shrinkage covariance estimator and proposed an explicit analytic estimator of it. The main result of the contribution is the proof of the convergence of the estimator under Kolmogorov asymptotics.
The numerical simulations highlight the gain in performance to add targets to the estimation. Experimentally, we argue that the convergence holds with only bounded fourth moments. The experiment in a non-stationary setup underlines the adaptability of our estimator in those situations, compared to other estimators.
Statements and Declarations
  • Funding: This research is funded by the CIFRE doctoral contract 2022/0536.
  • Conflict of interest/Competing interests: No conflict of interests nor competing interests.
  • Ethics approval and consent to participate: not applicable.
  • Consent for publication: all co-authors consent for publication.
  • Data availability: data is available online via yfinance API as explained in the corpus.
  • Materials availability: not applicable.
  • Code availability: code is available online via the GitHub linked in the corpus.
  • Author contribution: the author contributed to every stage of the article.
Appendix A Unknown mean formulas
We extend the bona fide estimator to setting of unknown mean, exposed in the introduction.